The article "Uncertainty Estimation in Railway Track Life-Cycle Cost"† presented the following data on time to repair (min) a rail break in the high rail on a curved track of a certain railway line.

159    120    480    149    270    547    340    43    228    202    240    218

A normal probability plot of the data shows a reasonably linear pattern, so it is plausible that the population distribution of repair time is at least approximately normal. The sample mean and standard deviation are 249.7 and 145.1, respectively.

(a) Is there compelling evidence for concluding that true average repair time exceeds 200 min? Carry out a test of hypotheses using a significance level of 0.05.
State the appropriate hypotheses.

H0: μ > 200
Ha: μ = 200H0: μ < 200
Ha: μ = 200    H0: μ = 200
Ha: μ ≠ 200H0: μ = 200
Ha: μ > 200H0: μ = 200
Ha: μ < 200


Calculate the test statistic and determine the P-value. (Round your test statistic to two decimal places and your P-value to four decimal places.)

t =
P-value =



What can you conclude?

There is compelling evidence that the true average repair time exceeds 200 min.

There is not compelling evidence that the true average repair time exceeds 200 min.    


(b) Using

σ = 150,

what is the type II error probability of the test used in (a) when true average repair time is actually 300 min? That is, what is β(300)? (Round your answer to two decimal places.)
β(300) =
Answer #1
Concepts and reason

The t-distribution is the continuous distribution where the sample size is small and population standard deviation is unknown.

If the test does not match with the reality then error occurred. There are 2 types of errors,

Type I and type II error.

Fundamentals

Null hypothesis states that there is no statistical significance exists in given set of observations, Type I error is rejecting H0{H_0} when H0{H_0} is actually true. The probability of type I error is represented by α\alpha .

Alternative hypothesis states that there is statistical significance exist in given set of observations. Type II error is accepting H0{H_0} when it is not true. The probability of type II error is represented by β\beta .

The probability value or p -value is the smallest value at which H0{H_0} can be rejected. It is used to determine the results of the hypothesis that is whether to reject or accept H0{H_0} . Power of the test is the probability of rejecting H0{H_0} when it is false. It is provided by 1β1 - \beta .

The test statistic of t-distribution is defined as,

t=XˉμS/n{\rm{ }}t = \frac{{\bar X - \mu }}{{S/\sqrt n }}

Where the sample mean is (Xˉ)\left( {\bar X} \right) , SS is sample standard deviation, nn is the sample size and μ\mu is the population mean. The p-value can be obtained by using excel function

=TDIST() = TDIST\left( {} \right)

Power of the test is defined as follows,

1β=P(Z>(zαμ1μ0σ/n))1 - \beta = P\left( {Z > \left( {{z_\alpha } - \frac{{{\mu _1} - {\mu _0}}}{{\sigma /\sqrt n }}} \right)} \right)

Where, zα{z_\alpha } is the critical value, σ\sigma is the population standard deviation, μ1{\mu _1} and μ0{\mu _0} are the two means.

(a.1)

The hypothesis to test the claim that true average repair time exceeds 200 is defined as,

The null and the alternate hypothesis can be stated as follows,

H0:μ=200H1:μ>200\begin{array}{l}\\{H_0}:\mu = 200\\\\{H_1}:\mu > 200\\\end{array}

So, the hypothesis Ha:μ=200,H0:μ<200{H_a}:\mu = 200,{H_0}:\mu < 200 , Ha:μ=200,H0:μ=200{H_a}:\mu = 200,{H_0}:\mu = 200 is incorrect as in alternative hypothesis true average repair time is equal to 200 which is not possible.

The hypothesis Ha:μ200,H0:μ=200{H_a}:\mu \ne 200,{H_0}:\mu = 200 is incorrect as the hypothesis is defined for two tail tests and to claim that true average repair time exceeds 200 is one tail test.

The hypothesis Ha:μ<200,H0:μ>200{H_a}:\mu < 200,{H_0}:\mu > 200 is incorrect as here the null hypothesis is incorrectly defined.

The hypothesis to test the claim that true average repair time exceeds 200 is defined as,

The null and the alternate hypothesis can be stated as follows,

H0:μ=200Ha:μ>200\begin{array}{l}\\{H_0}:\mu = 200\\\\{H_a}:\mu > 200\\\end{array}

(a.2)

The population distribution of repair time is approximately normal with sample mean as 249.7 and standard deviation as 145.1. The sample size is 12. So, the test statistic value of t-test is calculated as,

t=XˉμS/n=249.7200145.1/12=1.19\begin{array}{c}\\t = \frac{{\bar X - \mu }}{{S/\sqrt n }}\\\\ = \frac{{249.7 - 200}}{{145.1/\sqrt {12} }}\\\\ = 1.19\\\end{array}

The p-value is calculated using excel.

fx =TDIST(1.19,11,1)
DE
0.129548!

The p-value is 0.1295.

(b)

From the provided information,

μ1=300μ0=200σ=150n=12\begin{array}{l}\\{\mu _1} = 300\\\\{\mu _0} = 200\\\\\sigma = 150\\\\n = 12\\\end{array}

The zα{z_\alpha } can be calculated using NORMSINV()NORMSINV\left( {} \right) function of excel as follows,

--NORMSINV(0.05)
E
F
1.645

1β=P(Z>(zαμ1μ0σ/n))=P(Z>1.645300200150/12)=P(Z>0.6644)=P(Z<0.6644)\begin{array}{c}\\1 - \beta = P\left( {Z > \left( {{z_\alpha } - \frac{{{\mu _1} - {\mu _0}}}{{\sigma /\sqrt n }}} \right)} \right)\\\\ = P\left( {Z > 1.645 - \frac{{300 - 200}}{{150/\sqrt {12} }}} \right)\\\\ = P\left( {Z > - 0.6644} \right)\\\\ = P\left( {Z < 0.6644} \right)\\\end{array}

The P(Z<0.6644)P\left( {Z < 0.6644} \right) is calculated using the NORMSDIST()NORMSDIST\left( {} \right) function of excel as follows,

=NORMSDIST(0.6644)
0.75

Then,

1β=0.75β=10.75β=0.25\begin{array}{c}\\1 - \beta = 0.75\\\\\beta = 1 - 0.75\\\\\beta = 0.25\\\end{array}

Ans: Part a.1

The hypothesis to claim that true average repair time is,

H0:μ=200Ha:μ>200\begin{array}{l}\\{H_0}:\mu = 200\\\\{H_a}:\mu > 200\\\end{array}

Part a.2

The p-value and test statistic value is,

t

=

1.191.19

P-value

=

0.12950.1295

Part b

The value of probability of type II error or β\beta is 0.250.25 .